OpenStudy (anonymous):

How do I solve on the interval from [0, 2pi): 2cos^x-sinx-1=0 ??????? HELP PLEASE

OpenStudy (anonymous):

Write it out in latex please :)

OpenStudy (anonymous):

\[2\cos ^{2}x-\sin x-1=0\]

OpenStudy (saifoo.khan):

Use the identity: \[\sin^2x + \cos^2x =1\]

OpenStudy (anonymous):

I have been trying that, but I keep getting stuck

OpenStudy (saifoo.khan):

2(1-sin^2x) - sinx -1 =0 2 - 2sin^2x - sinx - 1 =0 -2sin^2x - sinx +1 =0 Now suppose sinx as y. Rewrite: -2y^2 - y + 1 = 0 Solve the quadratic.

OpenStudy (anonymous):

But I'm solving on an interval of 0 to 2pi, so it needs to be answers like pi/4 pi/6 etc...

OpenStudy (saifoo.khan):

Yes. Once you get the value of y, place them in the equation you made: sinx = y. Solve for y.

OpenStudy (anonymous):

My teacher never taught us to use y...so now I'm confused out of my mind...

OpenStudy (anonymous):

we're working with the unit circle

OpenStudy (saifoo.khan):

Alright then. Factor this out: -2sin^2x - sinx +1 =0

OpenStudy (anonymous):

would you factor out negative one?

OpenStudy (saifoo.khan):

You can do that too. Your answer will be same.

OpenStudy (mertsj):

\[-2\sin ^2x-\sin x+1=0\]

OpenStudy (mertsj):

\[2\sin ^2x+\sin x-1=0\]

OpenStudy (mertsj):

\[(2\sin x-1)(\sin x+1)=0\]

OpenStudy (mertsj):

Set each factor equal to 0 and solve.